OpenStudy (anonymous):

Find the x and y values for each of the item below a^xa^y=a^88 x and y difer by 32 a^x a^y=a^118

OpenStudy (anonymous):

First, you'll need to write an equation. Something like x+(x+32)=118 You then solve for x. x=43, y=43+32; 75. :D

OpenStudy (anonymous):

Can you explain the other one.

OpenStudy (anonymous):

What other one?

OpenStudy (anonymous):

a^xa^y=88

OpenStudy (anonymous):

Does it have anything like the y value is blahblahblah more than the other?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the x and y differ by thirty two relates to both problems, right? If so, same process. x+x+32=88. so 2x+32=88, isolate x, 2x=56 x=28, y=50